OpenStudy (anonymous):

A pile of gravel is shaped like a cone and has a radius of 7 ft and a height of 8 ft. Garden Landscaping Company purchases some of the gravel. The remaining pile has a radius of 4 ft and a height of 5 ft. How much gravel did Garden Landscaping Company purchase? Use 3.14 to approximate pi and round the answer to two decimal places.

OpenStudy (displayerror):

What you want to do is compute the volume of the two individual cones, then subtract the volumes in order to determine the volume of gravel that was purchased, \(V\). The volume of a cone is given by \[\text{Volume} = \frac{1}{3} \pi \left(\text{Radius}\right)^2 \times \text{Height}\] If you're looking for the mass, then look up the density of gravel (if there is one, I'm not sure) and use the following formula to convert from volume to mass: \[\text{Mass} = \text{Density} \times \text{Volume}\]

OpenStudy (anonymous):

Uhh, answer?